### v1

@@ 4.5. Evaluation Environments
5. Results and Analysis
5.1. FinSM performance
Table 5 presents the Hit-Rate@k performance of various LLMs on the FinSM task under zero-shot settings. The results show that LLMs still struggle to detect semantically inconsistent US-GAAP tags, with the best model (DeepSeek-V3) achieving only 11.89% average hit rate. This highlights the challenge of aligning models with fine-grained financial semantics. Several open-source models, including DeepSeek-V3, Qwen3-235B, and gemma-3-27b-it, outperform the closed-source GPT-4o, indicating that open-source progress is closing the gap in domain-specific retrieval. In contrast, financial domain models (Fin-o1-14B and Fin-R1) perform poorly, with Fin-o1-14B failing to produce any correct hits. This suggests a mismatch between their generative training focus and the retrieval-oriented nature of FinSM. Qwen3-32B also yields zero hit rates, reflecting possible instability in mid-scale variants. Overall, model scale contributes to performance gains (e.g., Qwen3-235B vs. Qwen3-32B; gemma-27B vs. gemma-12B), but the Llama series shows inconsistent scaling, signifying the importance of architecture and alignment strategies.
Table 5. Hit-Rate@k (%) performance under the zero-shot settings on the FinSM task.
LLMs $HR@1$ $HR@5$ $HR@10$ $HR@20$ $HR@AVE$
GPT-4o 7.88 9.09 9.09 9.09 8.79
DeepSeek-V3 11.52 11.82 11.82 12.42 11.89
Qwen3-235B-A22B-Instruct-2507 9.09 10.00 10.00 10.00 9.77
Llama-4-Scout-17B-16E-Instruct 1.82 2.42 2.73 3.03 2.50
Qwen2.5-72B-Instruct 7.88 8.18 8.48 8.48 8.26
Llama-3.3-70B-Instruct 5.15 5.15 5.15 5.15 5.15
Qwen3-32B 0.00 0.00 0.00 0.00 0.00
gemma-3-27b-it 10.00 10.30 10.30 10.30 10.23
gemma-3-12b-it 8.79 9.70 9.70 10.30 9.62
Llama-3.1-8B-Instruct 6.06 6.06 6.06 6.06 6.06
Llama-3.2-3B-Instruct 3.33 3.94 4.24 4.55 4.02
Fin-o1-14B 0.00 0.00 0.00 0.00 0.00
Fin-R1 2.12 2.12 2.42 2.73 2.35
Table 6 reports the Recall@k performance of different LLMs on the FinSM task under zero-shot settings. Overall recall values remain low, with the best model, DeepSeek-V3, achieving only 8.83% on average. This suggests that while some models can identify relevant tags, they fail to comprehensively retrieve all mismatched US-GAAP elements within the top-$k$ predictions. Large open-source models such as DeepSeek-V3, Qwen3-235B, and gemma-3-27b-it outperform GPT-4o, showing growing competitiveness in capturing broader semantic coverage. In contrast, financial domain models (Fin-o1-14B and Fin-R1) perform poorly, with Fin-o1-14B yielding zero recall, indicating a mismatch between domain-specific training objectives and retrieval-oriented evaluation. Similarly, Qwen3-32B fails to retrieve relevant tags, reflecting inconsistency in mid-scale variants. As with the Hit Rate results, scaling trends are non-monotonic: larger models often perform better, but the Llama series exhibits irregular improvements. These results highlight the difficulty of FinSM from a recall perspective, which demands that models capture all semantically inconsistent tags across interrelated XBRL filings and taxonomy documents.
Table 6. Recall@k (%) performance under the zero-shot settings on the FinSM task.
LLMs $R@1$ $R@5$ $R@10$ $R@20$ $R@AVE$
GPT-4o 4.93 6.82 6.98 7.01 6.43
DeepSeek-V3 7.06 8.82 9.33 10.11 8.83
Qwen3-235B-A22B-Instruct-2507 6.38 7.77 8.17 8.33 7.66
Llama-4-Scout-17B-16E-Instruct 0.92 1.58 1.91 2.21 1.66
Qwen2.5-72B-Instruct 4.31 4.81 5.18 5.18 4.87
Llama-3.3-70B-Instruct 3.34 3.78 3.85 3.85 3.71
Qwen3-32B 0.00 0.00 0.00 0.00 0.00
gemma-3-27b-it 5.91 7.81 8.20 8.26 7.54
gemma-3-12b-it 5.09 7.06 7.61 8.49 7.06
Llama-3.1-8B-Instruct 2.88 4.32 4.73 4.82 4.19
Llama-3.2-3B-Instruct 1.64 3.23 3.68 3.98 3.13
Fin-o1-14B 0.00 0.00 0.00 0.00 0.00
Fin-R1 1.28 2.02 2.32 2.63 2.06
Table 7 presents the Macro-F1@k results for the FinSM task under zero-shot settings. Compared with Hit Rate and Recall, Macro-F1 provides a stricter measure by jointly considering precision and recall, reflecting the balance between correctly retrieved mismatched tags and false positives. Overall scores remain very low, with the best model reaching only 9.35% on average, underscoring the difficulty of achieving both accuracy and coverage in structured financial retrieval. DeepSeek-V3 performs best across all cutoffs, followed by Qwen3-235B and gemma-3-27b-it, which also outperform GPT-4o, showing that open-source models are increasingly competitive in balancing precision and recall. In contrast, financial domain models perform poorly, with Fin-o1-14B producing no correct outputs and Fin-R1 reaching only 1.84%, suggesting that domain adaptation without retrieval alignment may increase false positives. Scaling effects appear within some model families, but the Llama series shows inconsistent improvements, indicating that architecture and alignment strategies influence performance more than model size. These findings highlight that FinSM requires models not only to retrieve relevant tags but also to control spurious predictions and maintain precision-recall balance across structured, multi-document inputs.
Table 7. Macro-F1@k (%) performance under the zero-shot settings on the FinSM task.
LLMs $MF1@1$ $MF1@5$ $MF1@10$ $MF1@20$ $MF1@AVE$
GPT-4o 5.43 6.91 7.03 7.02 6.60
DeepSeek-V3 7.73 9.54 9.98 10.17 9.35
Qwen3-235B-A22B-Instruct-2507 6.85 7.83 7.96 7.97 7.65
Llama-4-Scout-17B-16E-Instruct 1.08 1.55 1.76 1.80 1.55
Qwen2.5-72B-Instruct 4.73 5.29 5.42 5.42 5.21
Llama-3.3-70B-Instruct 3.65 4.07 4.14 4.14 4.00
Qwen3-32B 0.00 0.00 0.00 0.00 0.00
gemma-3-27b-it 6.47 8.38 8.67 8.71 8.06
gemma-3-12b-it 5.57 7.62 8.04 8.27 7.38
Llama-3.1-8B-Instruct 3.29 4.47 4.77 4.80 4.33
Llama-3.2-3B-Instruct 2.00 2.88 2.95 2.94 2.69
Fin-o1-14B 0.00 0.00 0.00 0.00 0.00
Fin-R1 1.46 1.92 1.97 2.00 1.84
Overall, the results confirm that FinSM presents a highly challenging setting based on structured, multi-document inputs. Across Hit Rate, Recall, and Macro-F1, current LLMs struggle with both coverage and the balance between precision and recall. Large open-source models such as DeepSeek-V3, Qwen3-235B, and gemma-3-27b-it outperform GPT-4o, whereas financial domain models fail to adapt, showing that domain specialization alone does not guarantee retrieval effectiveness. FinSM therefore serves as a rigorous benchmark for evaluating LLMs’ capacity for semantic retrieval and taxonomy alignment in realistic financial auditing contexts.
5.2. FinRE performance
Table 8 reports the overall performance of various LLMs on the FinRE task under zero-shot settings. The results demonstrate that relationship extraction is highly challenging for current LLMs, as it requires precise interpretation of hierarchical and compositional dependencies within structured XBRL filings and their alignment with the US-GAAP taxonomy. Among all models, GPT-4o achieves the best performance with 91.82% accuracy and balanced precision, recall, and F1 (around 90%), showing a strong capability to consistently identify different types of erroneous relationships. DeepSeek-V3 follows with 82.73% accuracy and 80.68% macro-F1, which, while competitive, indicates a noticeable gap compared to GPT-4o in balancing precision and recall.
@@ 5.2. FinRE performance
By contrast, the majority of open-source models lag significantly. Qwen2.5-72B and Qwen3-235B achieve moderate performance (Acc 67.50% and 62.73%, respectively), suggesting partial capability in capturing structural errors but with reduced robustness. Mid-scale models such as gemma-3-27b-it (Acc 45.91%) and gemma-3-12b-it (Acc 27.95%) perform weakly, while the entire Llama series struggles across scales. In particular, Llama-4-Scout-17B-16E-Instruct, despite being an extremely large-scale model, reaches only 27.50% accuracy and 22.71% macro-F1, performing worse than some mid-scale Qwen and Gemma models. This highlights that scale alone does not translate into structural reasoning ability without proper alignment to taxonomy-driven tasks. Notably, Qwen3-32B and the financial domain-specific Fin-o1-14B completely fail (0.00% across all metrics), while Fin-R1 remains weak with only 32.73% accuracy and 25.97% macro-F1, underscoring that domain specialization without explicit structural training does not generalize to relationship extraction.
Figure 3 reveals substantial variation across relation types. GPT-4o demonstrates the most consistent performance, exceeding 80% F1 on Reversal and Inappropriateness and achieving 100% on CombinationErr, showing strong robustness across all categories. DeepSeek-V3 follows with balanced results, though consistently lower than GPT-4o. In contrast, most open-source models struggle on CombinationErr, with many collapsing to near zero, indicating that axis-member validation across linkbases poses the greatest challenge. The Llama family performs poorly across all types, and both Qwen3-32B and Fin-o1-14B completely fail, achieving 0.00 in every category. Interestingly, Fin-R1 reaches a moderate 60.51 on CombinationErr but remains weak elsewhere. These results suggest that while top-tier general models can handle structural relationship extraction, most open-source and financial-domain models lack the structural alignment needed, with CombinationErr emerging as the hardest category overall.
@@
In summary, the FinRE results show that while frontier models such as GPT-4o achieve strong and balanced performance across all relation types, most open-source models lag considerably, and domain-specific financial LLMs fail to generalize to this structured classification setting. At the label level, CombinationErr emerges as the most difficult category, where even competitive models face a sharp drop, indicating that validating axis–member consistency requires reasoning across multiple interrelated documents. Overall, FinRE highlights that accurate financial relationship extraction demands more than scale or domain specialization, instead requiring explicit alignment with taxonomy-driven, cross-document reasoning.
5.3. FinMR performance
Figure 4. The accuracy (%) under the zero-shot settings on the FinMR task.
Figure 5. The error-rate results (%) for the FinMR task, where SER denotes the structural error rate, EER represents the extraction error rate, and CER refers to the calculation error rate.
Figure 4 reports the zero-shot accuracy on the FinMR task. Overall, all models show limited ability to perform structured mathematical reasoning, with the best model, Fin-o1-14B, reaching only 13.86% accuracy. This highlights the inherent difficulty of the task, which requires models to extract, compute, and verify numerical relations defined across hierarchical calculation structures in XBRL filings.
Across all LLMs, as shown in Figure 5, calculation errors (CER) dominate the performance gap. Most models exhibit extremely high CER values, typically between 70% and 83%, indicating that they struggle to follow multi-step arithmetic relationships or maintain consistency with domain-specific calculation rules. In contrast, extraction errors (EER) are comparatively moderate (roughly 10–22%), suggesting that LLMs can often locate the reported value but fail when performing reasoning-based computation.
Interestingly, Fin-o1-14B achieves the highest accuracy while maintaining a very low CER (9%), yet it suffers from a high structural error rate (SER = 71%). This pattern implies that once the model identifies a valid structural path, it performs reliable arithmetic reasoning but often fails to interpret the schema and linkbase structure correctly. Similarly, Qwen3-32B and Fin-R1 collapse due to severe structural errors, showing that without schema alignment, even accurate computation cannot be triggered.
Taken together, these results suggest that FinMR poses a fundamentally harder challenge than surface-level numerical QA. Success requires not only multi-step calculation ability but also schema-aware reasoning and structural grounding across multiple XBRL documents, capabilities that remain far beyond what current LLMs can reliably achieve.
@@ 6. Conclusion
In this work, we present 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark designed to evaluate LLMs on structured, hierarchical, and taxonomy-driven financial multi-documents. Built upon real-world XBRL filings, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: defines three complementary tasks, FinSM, FinRE, and FinMR, which jointly assess structured semantic consistency, hierarchical relationship understanding, and multi-step mathematical reasoning. Our extensive zero-shot evaluation of SOTA LLMs shows that even leading models such as GPT-4o, DeepSeek-V3, and Fin-o1-14B still struggle with cross-document reasoning, structured interpretation, and taxonomy alignment. These findings highlight the limitations of current LLMs in financial auditing and emphasize the importance of developing models with stronger structural grounding and reasoning awareness. We release all datasets and evaluation code to support future research on trustworthy financial intelligence.
@@ Appendix D Evaluation Models
  • Financial-specific LLMs: We include Fin-o1-14B (Qian et al., 2025) and Fin-R1 (Liu et al., 2025), both trained on domain-specific financial corpora to capture specialized terminology and reasoning patterns in financial reporting. These models allow us to assess the impact of domain adaptation on structured auditing and numerical reasoning tasks.

### v2

@@ 3.4. Evaluated LLMs
4. Benchmarking Results and Empirical Analysis
4.1. FinSM
Table 5. Zero-shot performance (%) on the FinSM task.
Hit Rate Recall Macro-F1
LLMs @5@5 @10@10 @20@20 @5@5 @10@10 @20@20 @5@5 @10@10 @20@20
GPT-4o 9.09 9.09 9.09 6.82 6.98 7.01 6.91 7.03 7.02
DeepSeek-V3 11.82 11.82 12.42 8.82 9.33 10.11 9.54 9.98 10.17
Qwen3-235B-A22B-Instruct-2507 10.00 10.00 10.00 7.77 8.17 8.33 7.83 7.96 7.97
Llama-4-Scout-17B-16E-Instruct 2.42 2.73 3.03 1.58 1.91 2.21 1.55 1.76 1.80
Qwen2.5-72B-Instruct 8.18 8.48 8.48 4.81 5.18 5.18 5.29 5.42 5.42
Llama-3.3-70B-Instruct 5.15 5.15 5.15 3.78 3.85 3.85 4.07 4.14 4.14
Qwen3-32B 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
gemma-3-27b-it 10.30 10.30 10.30 7.81 8.20 8.26 8.38 8.67 8.71
gemma-3-12b-it 9.70 9.70 10.30 7.06 7.61 8.49 7.62 8.04 8.27
Llama-3.1-8B-Instruct 6.06 6.06 6.06 4.32 4.73 4.82 4.47 4.77 4.80
Llama-3.2-3B-Instruct 3.94 4.24 4.55 3.23 3.68 3.98 2.88 2.95 2.94
Fin-o1-14B 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Fin-R1 2.12 2.42 2.73 2.02 2.32 2.63 1.92 1.97 2.00
From Table 5, we observe that retrieval performance on FinSM remains uniformly low across models, with no consistent advantage from larger parameter scales or domain-specific financial pretraining. Even state-of-the-art general-purpose LLMs achieve Hit Rate@20 below 13%, while several models fail. This suggests that FinSM performance is not primarily driven by model size or surface-level domain familiarity, but instead depends on the ability to identify semantically inappropriate accounting concepts under strict taxonomy-defined constraints.
Notably, increasing model capacity yields marginal improvements, indicating implicit representation learning is insufficient when hierarchical, taxonomy-defined semantics are not explicitly modeled. Financial models show similarly limited gains, highlighting that exposure to financial text does not guarantee accurate detection of mis-tagged concepts within regulated taxonomies.
The qualitative analysis reveals that retrieval errors arise from confusion over which reported concepts are semantically inappropriate under the given reporting context. For example, a filing incorrectly reports CashAndCashEquivalentsAtCarryingValue in a context where only Cash is permitted. In such cases, models often fail to flag the mismatch, instead retrieving tags or omitting the erroneous concept entirely. Similarly, models may overlook the misuse of aggregate tags such as AssetsCurrent when more granular components are required by the disclosure context. These errors indicate that models tend to rely on surface-level semantic similarity, rather than reasoning about contextual validity and taxonomy-defined reporting granularity.
4.2. FinRE
Table 6 indicates that FinRE is substantially more challenging than standard relation classification, as it requires precise interpretation of hierarchical, compositional, and cross-document constraints rather than surface-level semantic cues.
@@ 4.2. FinRE
Among all evaluated models, GPT-4o exhibits the strongest and most stable performance, achieving high accuracy and balanced macro-level precision, recall, and F1. This suggests that frontier models are capable of jointly reasoning over multiple relationship types when sufficient representational and reasoning capacity is available. DeepSeek-V3 follows as the strongest open-source model, but with a noticeable performance gap, indicating reduced robustness across relation categories.
In contrast, most open-source models struggle to generalize to this setting. Model scale alone does not lead to consistent improvements: several large models underperform smaller counterparts, and the Llama family exhibits uniformly weak results across scales. Notably, domain-specific financial LLMs do not demonstrate a clear advantage, with one model failing entirely and another achieving only limited accuracy. These observations suggest that relationship error classification in FinRE cannot be addressed by domain pretraining or parameter scaling alone, but instead requires explicit alignment with taxonomy-defined structural semantics.
@@
Finer-grained analysis by relation type (Figure 3) reveals substantial difficulty variation. While Reversal and Inappropriateness relations are handled reasonably well by top models, CombinationErr emerges as the most challenging category. Most models collapse to near-zero F1 on this relation type, indicating difficulty in validating axis–member combinations that require reasoning across multiple interrelated linkbases. Even competitive models exhibit sharp performance drops, highlighting the complexity of enforcing structural consistency beyond pairwise relationships.
4.3. FinMR
(a) Overall accuracy under zero-shot settings.
(b) Error-rate composition, including SER, ERR, and CER.
Figure 4. Zero-shot performance on the FinMR task.
Figure 4 summarizes the zero-shot performance of representative large language models on the FinMR task. Overall accuracy remains consistently low, with the best-performing model, Fin-o1-14B, achieving only 13.86% accuracy. This underscores the intrinsic difficulty of FinMR, which requires jointly reasoning over hierarchical calculation structures, cross-fact dependencies, and arithmetic constraints defined in XBRL filings, rather than answering isolated numerical questions.
The error composition analysis in Figure 4(b) provides a more fine-grained diagnostic view of model failures. Calculation errors (CER) dominate across most models, typically accounting for 70–83% of failures. This indicates that while models may identify relevant financial facts, they frequently fail to execute or validate multi-step arithmetic relations specified by taxonomy-grounded calculation rules. Extraction errors (EER) are comparatively moderate (approximately 10–27%), suggesting that locating candidate values is often feasible, but transforming them into correct numerical inferences remains a major bottleneck.
Structural errors (SER) further expose limitations in producing structure-compliant outputs. Models such as Qwen3-32B and Fin-R1 exhibit extremely high SER values, indicating frequent failures to generate outputs that conform to the required structured format. In contrast, Fin-o1-14B shows a markedly different error profile: despite achieving the highest overall accuracy and maintaining a low CER (9%), it still suffers from a high SER (71%). This pattern suggests that even when numerical reasoning is reliable, failures in adhering to the expected output format can prevent correct answers from being properly expressed and evaluated.
@@ 5. Conclusion
In this work, we present 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark designed to evaluate LLMs on structured, hierarchical, and taxonomy-driven financial reasoning across multi-document filings. Built from real-world XBRL data, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: defines three complementary tasks, FinSM, FinRE, and FinMR, which jointly assess structured semantic consistency, hierarchical relationship understanding, and multi-step mathematical reasoning. Our extensive zero-shot evaluation of state-of-the-art LLMs shows that even leading models, such as GPT-4o, DeepSeek-V3, and Fin-o1-14B, still struggle with cross-document reasoning, structured interpretation, and taxonomy alignment. These findings highlight fundamental limitations in current LLMs for financial auditing and emphasize the importance of developing models with stronger structural grounding and explicit reasoning awareness. We release all datasets and evaluation code to support future research on trustworthy financial intelligence.
@@ Appendix D Evaluation Models
  • Financial-specific LLMs: We include Fin-o1-14B (Qian et al., 2025b) and Fin-R1 (Liu et al., 2025), both trained on domain-specific financial corpora to capture specialized terminology and reasoning patterns in financial reporting. These models allow us to assess the impact of domain adaptation on structured auditing and numerical reasoning tasks.
@@
Appendix E Practical Deployment and Scalability Considerations
This section provides additional discussion on how FINAUDITING may be integrated into real-world auditing workflows and scaled under practical constraints. The goal is not to prescribe a specific deployment system, but to illustrate how the benchmark design aligns with existing auditing practices and supports modular instantiation.
Integration into Existing Auditing Pipelines
In practice, financial auditing commonly relies on rule-based validation systems, such as XBRL consistency checks and Data Quality Committee (DQC) rules, followed by targeted human review. FINAUDITING is naturally positioned to operate downstream of such validators. Rather than processing entire filings, models evaluated under FINAUDITING are applied only to a subset of suspicious items already identified by deterministic checks. Within this setting, the three benchmark tasks correspond to distinct stages of post-validation analysis. FinSM supports taxonomy disambiguation by assisting auditors in identifying appropriate concepts when potential mis-tagging is detected. FinRE provides structured categorization of detected issues into interpretable error types, which can guide remediation and prioritization. FinMR focuses on targeted numerical verification by tracing calculation structures and comparing reported values with recomputed ones, and is intended for precision-critical cases. This workflow reflects common auditing practices in which automated components emphasize recall, while high-stakes decisions remain subject to expert review.
Scalability across Regulatory Environments
Although instantiated using the US-GAAP taxonomy, the task formulation of FINAUDITING is not tied to any specific accounting standard. The benchmark abstracts auditing as taxonomy-grounded reasoning over instance documents, schema definitions, and linkbase relations. Extending FINAUDITING to other regulatory environments, such as IFRS or jurisdiction-specific reporting standards, primarily requires substituting the underlying taxonomy and rule sources while preserving the same task interfaces. This decoupling between task definitions and accounting standards enables scalable adaptation without redesigning the benchmark structure.
Lightweight Variants and Semi-automatic Deployment
FINAUDITING also supports tiered or hybrid deployment strategies that balance performance with computational cost. Full-capacity large language models are not required at all auditing stages. For example, FinSM can be instantiated using embedding-based retrieval or compact instruction-tuned models to achieve high recall at low cost, effectively narrowing the candidate space for auditors. FinRE involves a small and fixed label set, making it suitable for fine-tuning lightweight classifiers. FinMR, which requires precise numerical reasoning and evidence tracing, can be selectively applied to a limited number of high-risk cases and handled by larger models or human-in-the-loop verification. Such semi-automatic configurations enable practical deployment by combining lightweight automation with selective use of more advanced reasoning components.
Overall, by aligning with existing auditing workflows and supporting model stratification, FINAUDITING provides a modular evaluation framework that reflects how machine-assisted financial auditing systems may be integrated and scaled under realistic operational constraints.